The difference between two positive numbers is 16. If the smaller of those two numbers is 3/5th of the larger, what is the value of the smaller number?
Answer by oberobic(2304)

x - y = 16
so
x = 16+y
or
y = x-16
.
y = 3/5*x
.
y=y
so
x-16 = 3/5*x
.
multiply both sides by 5 to eliminate the fraction
.
5(x-16) = 5(3/5*x)
.
5x -80 = 3x
.
2x = 80
.
x = 40
.
y = 40-16
y = 24
.
Is 24 3/5 of 40?
3*40 = 120
120/5 = 24
Yes, it is.
.
Answer: The value of the smaller number is 24.
.
Done.